FINDING THE ELUSIVE IODOCARBENE: FLUORESCENCE EXCITATION AND SINGLE VIBRONIC LEVEL EMISSION SPECTROSCOPY OF CHI

C. TAO, C. EBBEN, H. T. KO AND \underline{S. A. REID}, Department of Chemistry, Marquette University, Milwaukee, WI 53233; Z. WANG AND T. J. SEARS, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973.

Among the triatomic halocarbenes, only the iodocarbenes remain to be characterized. The search for these elusive species is motivated by a controversy regarding the multiplicity of the ground state. Photoelectron spectra of Lineberger and co-workers ^{a} suggest a triplet ground state for CHI, at variance with recent ab initio studies, which suggest a singlet ground state with a singlet-triplet gap of around 4 kcal mol^{−1}. In this work, we have succeeded in finding the spectra of CHI and its deuterated isotopomer using pulsed discharge jet spectroscopy. Rotationally resolved fluorescence excitation spectra are consistent with a singlet-singlet transition, and the derived rotational constants are in good agreement with theoretical predictions. Single vibronic level emission spectra confirm a singlet multiplicity for the ground state, and reveal extensive mixing of the singlet and triplet levels at higher energy. We are able to set a lower limit on the singlet-triplet gap of 4.1 kcal mol^{−1}, in excellent agreement with theory. Extrapolation of the observed bending levels for CHI and CDI to a common origin suggests that the origin of the A^{1}A′′ state lies near 10 500 cm^{−1}, and we will report on high resolution measurements near the electronic origin made at Brookhaven National Laboratory.

aM. K. Gilles, K. M. Ervin, J. Ho, and W. C. Lineberger, J. Phys. Chem. 96, 1130 (1992).